Three point mapping:

A, B, and C genes (alleles A and a, B and b, C and c) with complete dominance for all.

The testcross is:     AaBbCc x aabbcc

The progeny are seen below.

ABC: 75          aBC: 75

ABc: 5             aBc: 345

AbC: 345        abC: 5

Abc: 75           abc: 75

1) If the genes are not linked, could this happen by chance alone? Write your assumptions. do the chi-square/show the math, and do you accept or reject the hypothesis that the genes are not linked and assort independently of each other?

2) Assume the genes are linked. Establish the Gene Order and map distances for the 3 loci, and draw a final map.
Answer #1

Answer:

1).

Phenotype

Observed(O)

Expected (E)

O-E

(O-E)2

(O-E)2/E

ABC

75

125

-50

2500.00

20.00

ABc

5

125

-120

14400.00

115.20

AbC

345

125

220

48400.00

387.20

Abc

75

125

-50

2500.00

20.00

aBC

75

125

-50

2500.00

20.00

aBc

345

125

220

48400.00

387.20

abC

5

125

-120

14400.00

115.20

abc

75

125

-50

2500.00

20.00

Total

1000

1000

1084.80

Chi-square value = 1084.80

Degrees of freedom = number of phenotypes – 1

Df = 8-1 = 7

The chi-square value of 1084.80 is greater than the critical value of 14.07 (at 5% p value). Hence, we reject null hypothesis and the genes are not assorted independently. The three genes are linked.

2).

Answer:

Hint: Non-recombinant (parental) combinations are produced more in number than recombinants. The parental (non-recombinant) genotypes is AbC / aBc

1).

If single crossover occurs between a & b

Normal combination: Ab/aB

After crossover: AB/ab

AB progeny=75+5=80

ab progeny = 5+75=80

Total this progeny = 160

Total progeny = 1000

The recombination frequency between a&b = (number of recombinants/Total progeny) 100

RF = (160/1000)100 = 16%

2).

If single crossover occurs between b&c

Normal combination: bC/Bc

After crossover: bc/BC

bc progeny= 75+75=150

BC progeny = 75+75=150

Total this progeny = 300

The recombination frequency between b&c = (number of recombinants/Total progeny) 100

RF = (300/1000)100 = 30%

3).

If single crossover occurs between a & c

Normal combination: AC/ac

After crossover: Ac/aC

Ac progeny=5+75=80

aC progeny = 5+75=80

Total this progeny = 160

Total progeny = 1000

The recombination frequency between a&c = (number of recombinants/Total progeny) 100

RF = (160/1000)100 = 16%

Recombination frequency (%) = Distance between the genes (mu)

b--------16mu------a----16mu----c

The order of genes = bac

Fingal gene map = b--------16mu------a----16mu----c